Find the determinant of the matrix. Expand by cofactors [ 7-11 1-5 10 (a) Row 2 (b) Column 2 284 noints SOLICA
Answer #1

RC21=(-1)2+(-1)= 1 RC22 = (-1)2+2 (7)=7 Determinant is = RC99(10) +RC2+(-5) =(7)(10)+(1)(-5) = 70-5 = 65 - - - - - CC2 = (-1)